OpenStudy (secret-ninja):

Am I right? Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 -5|3x - 2| = -15 |2x - 9| = 11 A. Equation #1 and Equation #2 have the same number of solutions. B. Equation #1 has more solutions than equation #2. C. Equation #1 has fewer solutions than equation #2. D. None of the statements above describe the number of solutions to equations shown.

OpenStudy (secret-ninja):

I think its A.

OpenStudy (anonymous):

I believe you're right. Moduli almost always give 2 distinct answers. You may want to solve and double check to be sure though.

OpenStudy (secret-ninja):

I did solve. For the first equation I got x=-2.6(6 is repeated) and x=-4 Fot the second equation I got x=10 and x=1

OpenStudy (radar):

Looking at Equation 2, it appears it would have two solutions.

OpenStudy (anonymous):

Yup, so A would be the answer.

OpenStudy (secret-ninja):

Thanks :]

OpenStudy (anonymous):

Anytime, if you have any more queries, don't hesitate to ask!

OpenStudy (secret-ninja):

Okie, Imma give you a medal swaschan, and you give a medal to radar so we are all even. :] You both helped ^.^

OpenStudy (radar):

2x - 9 = 11 2x - 9 = -11

OpenStudy (radar):

Solving 2 would give you x = 10 and x = -1. While looking at equation it looks like it may be a little different as the absolute value is then multiplied by -5

OpenStudy (radar):

-5|3x - 2| = -15 |3x-2| =-15/-5 = 3 I guess we would now solve same as before. 3x - 2 = 3 and 3x -2 = -3 ist cond. 3x = 5, x = 1 2/3 2nd cond 3x = -1 giving x = -1/3 Should verify by substituting in the original equation. -5|3(5/3) - 2)| = -15 -5 (3) = -15 yup -15 = -15 other solutions check -5|3(-1/3) - 2| = -15 -5(3) = -15, yup -15 = -15 Just had to prove it to myself, agree A

OpenStudy (radar):

Should of been most obvious to the casual observer lol. Still proof is good.